The cardiologist orders Ralph to receive heparin according to the following: 80 units/kg IVP once (this...

  1.      The cardiologist orders Ralph to receive heparin according to the following: 80 units/kg IVP once (this is called a bolus dose), then 18 units/kg/hr infusion.  Ralph weighs 231 pounds.  You have the following available to you: heparin 20,000 units in 250 mL D5W and 1-mL vials of heparin that is 5000 units/mL.  How many mL will you bolus Ralph with?  You’ll then set the infusion up on a pump; how many mL per hour will you set the pump rate?
Answer #1

1. 231 pounds = 104.78 kg

Total heparin needed for bolus dose = 80× 104.78 ie, 8382.4 units

1ml vial =   5000 units heparin

Then ralph needs ( 8382.4 ÷ 5000) = 1.6 ml

Bolus dose .

2. Step 1.

Starting units per hour =18 units ×104.78 kg ie, 1886.04 units /hour.

Step 2.

Calculate the rate of infusion per hour

20000 units ÷ 250 ml =1886.04 units /hour ÷ X (ml/hr )

20000 units × (X ml/hr ) = 471510

X ml/hr = 471510 ÷ 20000

X = 23.5 ml/ hr

Hence 23.5 ml/hr is the heparin infusion rate.